# LASLiN: A Learning-Augmented Peer-to-Peer Network

### v1

@@ 6 Learning-Augmented Continuous SLN
In this section, we present LASLiN, a learning-augmented Continuous SLN protocol where each node can obtain a prediction of the future demand matrix at any time.
@@
We first show that, for general demands, any randomized, demand-oblivious P2P protocol can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN. This result provides a comparison point around which the consistency and robustness metrics of demand-aware algorithms can be evaluated.
Theorem 6.1.
@@
We consider any randomized P2P protocol (Definition 5.1) that is ideal (Definition 2.4) and not learning-augmented (Definition 2.7). Then, there exists a demand matrix $W$ such that the considered protocol has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN protocol pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
Proof 6.2.
@@
As the considered protocol is ideal, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN protocol can clearly pay a cost of $1$, the considered randomized SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
@@
Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$. We considered all the cases and the claim follows.
Protocol 2 (LASLiN).
@@
LASLiN is a randomized, continuous SLN that sets its heights by adding the input, integer prediction in $1,\lceil\ln n\rceil]$, and a real number in $(0,1)$ chosen uniformly at random.
Theorem 6.3.
LASLiN is a randomized, Continuous SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust.
Proof 6.4.
We start by proving that the maximum degree in LASLiN is lower than $4\ln^{2}n$ with high probability. Let $u$ be a node, let $p_{u}\in[1,\lceil\ln n\rceil]$ be the predicted height of $u$ and let $r_{u}\in(0,1)$ be the uniform random noise added on top of $p_{u}$. We partition the neighbors of $u$ into $p_{u}$ subsets called $S(1),S(2)\dots S(p_{u}+1)$: For all $i\in[1,p_{u}]$, $S(i)$ contains the neighbors with height between $i$ and $i+1$, the last set $S(p_{u}+1)$ contains the neighbors with height greater than $p_{u}+1$. According to Theorem 5.8, each set from $S(1)$ to $S(p_{u})$ contains fewer than $4\ln n$ nodes with probability greater than $1-\frac{1}{n^{4}}$ (this is the result before the last paragraph in the proof) — the set $S(p_{u}+1)$ contains at most two nodes. We conclude with a union bound that $u$ has fewer than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{3}}$, and that the maximum degree is lower than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{2}}$ also with a union bound.
@@
The $O(1)$-consistency is direct as the comparative optimal static SLN has its heights within a $O(\log n)$ factor without loss of generality as its maximum degree is bounded by $O(\log n)$ by assumption. We now prove the robustness claim. A routing path between any two nodes $u$ and $v$ has two phases: first ascending (the heights always increase) and then descending (the heights always decrease). For both phase, the number of different encountered integer heights is upper-bounded by $O(\log n)$. For each encountered integer height, the expected routing path length is upper-bounded by $O(\log n)$ (Theorem 5.2). We finish the proof with an upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
We now define the notion of super-sparse matrix, which indicates a concentration of the weights of the matrix on a very small (logarithmic) number of coefficients.
Definition 6.5 (super-sparse matrix).
A matrix $W\in M_{n}(\mathbb{R})$ is super-sparse if there exists a constant $c\geq 1$ and a restricted subset of node pairs $S\subseteq V^{2}$ such that:
$|S|$ $\leq O(\log^{c}n)\qquad\text{ and }$
$\sum_{(u,v)\in S}W(u,v)$ $\geq\left(1-O(1/\log n)\right)\cdot|W|$
Theorem 6.6.
Assume that the demand matrix $W$ is super-sparse. We consider the protocol Binary LASLiN (B-LASLiN) where the input prediction $p$ can only have two values, $1$ or $2$. B-LASLiN is $O(\log\log n)$-consistent and $O(\log n)$-robust.
In asymptotics, the performances are therefore always better than any demand-oblivious protocol, with an exponential improvement in case of perfect predictions.
Proof 6.7.
We start by proving that the maximum degree of each node is lower than $O(\log n)$ with high probability. As seen in the proof of Theorem 6.3, each node $u$ has fewer than $p_{u}\cdot\ln n$ neighbors with high probability, the claim follows since $p_{u}\leq 2$.
We prove that the expected total weighted path length is lower than $O(\log n)\cdot|W|$ regardless of the input predictions. The proof is similar to the proof of Theorem 6.3 except, as the predictions are upper-bounded by a constant, we avoid the additional $\log n$ factor.
Finally, we prove that the expected total weighted path length is lower than $O(\log\log n)\cdot|W|$. We consider the following set of predictions: let $S$ be the node pairs from Definition 6.5, we give prediction $2$ to the nodes that have at least one coefficient from $S$ in their row or column in $W$, otherwise $1$. We call $V_{2}$ ($V_{1}$) the set of nodes that have prediction 22 (11). As $W$ is super sparse, it holds that $|V_{2}|\leq O(\log^{c}n)$.
$\sum_{u,v}W(u,v)\cdot d(u,v)$ $=\sum_{(u,v)\in S}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq\sum_{u,v\in V_{2}}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq O(\log\log n)\cdot|W|+O\left(\frac{1}{\log n}\right)\cdot|W|\cdot\ln n$
$=O(\log\log n)\cdot|W|$
To go from line 3 to 4 we used (1) Theorem 5.2 on the routing paths between the nodes of $V_{2}$ — there are $O(polylog(n))$ of those nodes, they have the largest heights of the SLN, so a path has expected length $\ln O(polylog(n))=O(\log\log n)$ — and on the routing paths from nodes in $V^{2}\setminus S$, and (2) we used Definition 6.5 to upper-bound the sum of coefficients in $W$ outside of $S$.
Figure 6: Comparison of the costs of multiple P2P networks against different demand matrix distributions. The demand matrices are generated based on the Zipfian distribution with varying skewness parameter $\zeta\in\{1.1,1.2\dots 2.0\}$, each coefficient is drawn independently and identically distributed. For each demand matrix distribution, we compare the cost of three P2P networks: the optimal static SLN, LASLiN, and the demand-oblivious Uniform SLN.
@@ 7 Experimental Evaluation
Demand matrices
In our experiments, we generate communication demand based on the Zipfian distribution [1], which receives a parameter $\zeta>0$, that indicates the distribution’s skewness. Roughly speaking, lower values of $\zeta$ result in a sparser demand matrix. The number of nodes in all experiments was set to $n=32$, and each experiment was repeated $30$ times.
Noise generation
As common in the literature [20], we study the robustness of our algorithms by introducing noise in the predicted data. The most used type of noise is additive, often gaussian [26] or lognormal [20]. This type of noise however almost did not impact the performance of our algorithms. Instead, we test two other models: the demand error recomputes a portion (x-axis) of the coefficients with the same Zipfian distribution, and the unit heights enforce that each node can receive a predicted height of $1$ with a certain probability (x-axis).
Results
In Figure 6 we evaluate the consistency of LASLiN in scenarios where the demand follows the Zipfian distribution, with varying parameter $\zeta\in[1.1,2]$, and where the prediction advice is perfect (i.e., there is zero noise in the demand matrix). We compare the total communication cost (as defined in (2.5)) of LASLiN to that of the static optimum SLN and the demand-oblivious SLN (Uniform) baselines. We can see that, in contrast to the demand-oblivious SLN, LASLiN obtains close-to-optimum total communication cost in all workloads, regardless of their spatial locality, i.e., LASLiN is consistent.
In Figure 7 we evaluate the robustness of LASLiN in scenarios with increasing demand noise (x-axis). We compare the total cost of LASLiN to that of the static optimum SLN, the demand-oblivious (Uniform) SLN, and the SLN baselines. We can observe that the SLN performs increasingly worse (has higher cost) the greater the amount of change (noise) in the input demand matrix. LASLiN, on the other hand, achieves consistently lower communication cost (than the former), regardless of the type of noise and regardless of the prediction fault probability, i.e., LASLiN is robust relative to both type and magnitude of prediction errors.
Figure 7: Cost of different P2P networks under varying noise amplitudes on the demand matrix predictions. Simulation results: robustness ($\zeta=1.2$)
@@ 8 Related Work
Improved designs of P2P networks. P2P networks have been the focus of computer scientists for a long while (see [19] for a detailed survey). In terms of P2P networks with constant degree, the notable contribution of [23] that extends the work of [13], giving a general P2P construction such that for a network with degree $d$ and $n$ nodes, you can have guaranteed path length of $\Theta(\log_{d}n)$. Lately, improving the design of networks given prior information about communication demand has grabbed the attention of researchers. In particular, the work of [7] set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking [11, 24, 12].
@@
Works closest to ours are by [6, 10] in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input. This is in particular important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system. Also, the work of [21] discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide any theoretical upper bound for the communication time. Lastly, we point out to other extensions of skip list, namely Skip Graph [3] and SkipNet [16] that, to facilitate P2P communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
@@
Learning-augmented algorithms. Given the rise of machine learning algorithms (that are prone to inherent unbounded failures), and after the pioneering work of [26] on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice. Such an approach has been used in various applications so far222see a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching [20], online TSP [15], list update [8].
@@
Some of the more recent works focus on improving single-source search data structures. Initially, the works of [18, 9] provided learning-augmented search trees. The works that are closest to our work are by [14, 30], which studied learning-augmented single-source skip lists. In contrast, we consider a more challenging model, where the focus is on pairwise communications rather than communication from a single source. Recently, we became aware of a not peer-reviewed manuscript [2] that aims to augment P2P networks with learned advice; however, this paper looks into improving self-stabilization and recovery of P2P networks, not improving communication delay.
9 Conclusion
In this work, introduced a continuous variant of Skip List Network (SLN), Continuous SLN. This variant is arguably easier to manipulate and analyze. In addition we show that we can use it to reach strong robustness properties even in an adversarial, fully distributed setting. Lastly, through experimental evaluation, we observed that our suggested algorithms outperform the theoretical results. As a future work, one can think of an alternative algorithm for OSSLN with a lower computational complexity.

### v2

@@ 5 Learning-Augmented C-SLN
In this section, we present LASLiN, a learning-augmented C-SLN algorithm, where each node can obtain a prediction of the future demand matrix at any time.
@@
We first show that, for general demands, any demand-oblivious overlay algorithm can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN. This result provides a comparison point around which the consistency and robustness metrics of demand-aware algorithms can be evaluated.
Theorem 7.
@@
We consider any overlay algorithm that is not learning-augmented (6). Then, there exists a demand matrix $W$ such that the considered algorithm has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN algorithm pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
Proof.
@@
As the considered algorithms has a logarithmic maximum degree with high probability, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN algorithm can clearly pay a cost of $1$, the considered SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
@@
Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$. We considered all the cases and the claim follows. ∎
Algorithm 2 (LASLiN).
@@
LASLiN is a learning-augmented C-SLN that takes an integer in $1,\lceil\ln n\rceil]$ as a prediction. Its initialization procedure set its height by adding the input prediction with a real number in ]0,1[]0,1[ chosen uniformly at random.
Theorem 8.
LASLiN is a C-SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust, while its maximum degree is $O(\log^{2}n)$ with high probability.
Proof.
We start by proving $O(1)$-consistency. Adding values in ]0,1[]0,1[ to the node’s heights does not remove the edges of the initial SLN. Let $u$ and $v$ be two nodes and consider the routing path $P_{u,v}=u,x_{1},x_{2}\dots v$ from $u$ to $v$ in the initial SLN. Now calling $P_{u,v}^{\prime}$ the path from $u$ to $v$ in the newly created C-SLN, we have that $P_{u,v}^{\prime}\subseteq P_{u,v}$; this can be show with a simple induction on the path length starting from $u$. All paths of the C-SLN have equal or shorted length than in the initial SLN with perfect predictions and $O(1)$-consistency follows.
@@
We now prove the robustness claim. A routing path between any two nodes $u$ and $v$ has two phases: first ascending (the heights always increase) and then descending (the heights always decrease). For both phase in the initial SLN, the number of different encountered integer heights is upper-bounded by $O(\log n)$. For each encountered integer height, the expected routing path length in the C-SLN is upper-bounded by $O(\log n)$ (Theorem 3) with high probability, we deduce that the routing path is lower than $O(\log^{2}n)$ with high probability using a union bound on the integer heights of the initial SLN. We now use this upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
Finally, we show the claim on the maximum degree. Let $u$ be a node. Fixing an integer $i\in[0,p_{u}]$, in the newly created C-SLN $u$ has fewer than $\ln n$ neighbors with initial integer heights $i$ with high probability (Theorem 6). We complete the proof with a union bound on all possible $O(\log n)$ heights of the initial SLN. ∎
Figure 3: Consistency (Pareto $\alpha\in[0.2,2]$, $n=64$, $\eta=0$).
(a) $\alpha=0.2$
(b) $\alpha=0.4$
(c) $\alpha=0.6$
(d) $\alpha=0.8$
(e) $\alpha=1.0$
(f) $\alpha=1.2$
(g) $\alpha=1.4$
(h) $\alpha=1.6$
(i) $\alpha=1.8$
(j) $\alpha=2.0$
Figure 4: Smoothness against Stale Noise (Pareto demand, $n=64$): average communication cost relative to uOPT.
(a) $\alpha=0.2$
(b) $\alpha=0.4$
(c) $\alpha=0.6$
(d) $\alpha=0.8$
(e) $\alpha=1.0$
(f) $\alpha=1.2$
(g) $\alpha=1.4$
(h) $\alpha=1.6$
(i) $\alpha=1.8$
(j) $\alpha=2.0$
Figure 5: Smoothness against Adversarial Noise (Pareto demand, $n=64$): average communication cost, relative to uOPT.
(a) pFabric traces ($\eta=0$)
(b) pFabric trace_0_1\_0\_1
(c) pFabric trace_0_5\_0\_5
(d) pFabric trace_0_8\_0\_8
Figure 6: pFabric demand under Adversarial Noise ($n=144$): consistency (a) and smoothness (b),(c),(d).
(a) Pareto $\alpha=0.2,n=64$
(b) pFabric trace_0_5,n=144\_0\_5,n=144
Figure 7: Node degree CDF ($\eta=0$): the vertical lines indicate the average node degree.
@@ 6 Experimental Evaluation
In this section, we present our experimental setup and results.
Algorithms and baselines.
To evaluate the performance of LASLiN, we implemented the following algorithms:
  • Sln: Demand-Oblivious Discrete SLN (1, with the height of each node having been assigned the number of consecutive independent trials in which a random variable $x\in[0,1]$ satisfied $x<0.5$);
  • uOPT: Unbounded Optimum Static SLN (4) on predicted demand;
  • bOPT: Bounded Optimum Static SLN (8 on predicted demand, with $h_{\max}=6$);
  • bLasin: LASLiN (6, with bOPT as input prediction);
Demand matrices.
In our experiments, we used both synthetic and real-world workloads. The synthetic workloads were based on a Pareto distribution, with the shape parameter $\alpha\in[0.2,2.0]$ (the higher the value of $\alpha$, the closer the demand to a uniform distribution). For each value of $\alpha$, we generated $30$ distinct demand matrices on a network with $64$ nodes. To validate our algorithms on real-world data, we utilized three pFabric traces Alizadeh et al. (2013), generated by executing simulation scripts in NS2 on a network with $144$ nodes.
Noise generation.
In line with some previous related work  Lykouris and Vassilvitskii (2018), we evaluate the smoothness of our algorithms by introducing noise, or prediction error, either to the demand weight values (between pairs of nodes) or to the heights assigned to nodes by an SLN algorithm. In particular, we implemented two noise models:
  • Stale Noise: This model aims to simulate the effect of an outdated input prediction through a probabilistic swapping mechanism. For every node pair in the demand matrix, the demand weight value is replaced by a different value, taken from a historical or reference workload, with probability $\eta$ (the noise intensity parameter).
  • Adversarial Noise: This model simulates the effect of the presence of malicious nodes in the network. For every network node, the height value assigned to it by the Skip List algorithm is reset to $1$, with probability $\eta$.
The Stale Noise model represents scenarios with delayed updates of input prediction values, caused by, e.g., network reconfigurations or synchronization errors. We restricted the evaluation of smoothness against stale noise to demand matrices based on Pareto distributions. [^1] We excluded pFabric traces from this part of the evaluation, as the dataset lacked sufficient independent traces to generate reliable reference models. Given a workload $W_{i},1\leq i<30$, we designated the subsequent workload, $W_{i+1}$, as the reference source for the current workload, $W_{i}$. Specifically, for each communication pair $(x,y)$, the actual demand weight $W_{i}(x,y)$ was replaced by the reference weight $W_{i+1}(x,y)$, with probability $\eta$.
In the Adversarial Noise model, adversarial agents reset their own node heights to a baseline value of $1$, with probability $\eta$, eliminating structural (SLN) shortcuts and increasing the total routing work (communication cost). The objective of this experiment is to demonstrate the efficacy of a simple countermeasure: by increasing each node’s height value with a small randomized value, we can effectively safeguard the network’s performance against such adversarial intent.
Metrics.
We used the following performance metrics:
  • Communication cost (3);
  • Consistency (7) and smoothness;
  • Node degree cumulative distribution function (CDF): percentage of nodes with degree $\leq x,x\in[1,n)$, and the average node degree.
Results.
In Figures 3 and 6(a), we evaluate the consistency of LASLiN under Pareto-distributed demand with variable shape parameter ($\alpha\in[0.2,2.0]$), assuming perfect prediction advice ($\eta=0$). We compare the (average) total communication cost of LASLiN (bLaslin) against three baselines: the bounded static optimum (bOPT), the unbounded static optimum (uOPT), and the demand-oblivious discrete SLN (Sln). The results demonstrate that, unlike the demand-oblivious baseline, LASLiN achieves a total communication cost close to the optimum across all simulated workloads, regardless of spatial locality ($\alpha$). This empirically confirms that LASLiN is consistent.
Notably, LASLiN, despite being projected onto the bOPT structure (bLaslin), i.e., using the solution of bOPT as its input prediction, actually achieves better communication cost than bOPT, even in noiseless scenarios (Fig. 3). This performance gain is enabled by the slight increase in the network’s average and maximum degrees, as shown in Figures 7(a) and 7(b). This behavior has been expected, since bLaslin effectively breaks height ties, inherent in the bOPT solution. By adding a small random remainder to the predicted heights, nodes that have previously shared the same rank (height) become differentiated, generating additional (SLN) shortcut links that further reduce routing cost.
In Figure 4, we evaluate the smoothness of LASLiN against increasing levels of stale noise. Both uOPT and bOPT are significantly impacted by the intensity of the stale noise (increasing $\eta$ on the x-axis), especially for lower $\alpha$ values, where the workload is more skewed. In contrast, bLaslin consistently achieves lower communication cost than its input prediction (bOPT). This demonstrates that LASLiN remains robust relative to both the type and magnitude of input prediction errors.
Additionally, we observe that, as $\alpha$ increases, approaching a uniform demand distribution, the performance of the optimum SLN converges to that of the discrete demand-oblivious SLN.
The strongest advantage of LASLiN is observed in the presence of adversarial noise (Figures 5 and 6(b,c,d), note that the y-axis of these plots is shown in the log scale). By simply projecting the LASLiN mechanism onto the predicted node heights, we successfully mitigate the disruption that could be caused by adversarial agents. We observe that while both uOPT and bOPT degrade progressively as noise intensity increases, eventually underperforming even the demand-oblivious discrete SLN baseline, bLaslin consistently achieves lower communication costs than both its reference prediction (bOPT) and the lower bound baseline (uOPT).
Finally, it is important to note that the average and the maximum node degree values obtained in the simulations (Fig. 7) are in alignment with the theoretical expectation (Theorem 5, Theorem 6 ). In particular, for the discrete SLN, the maximum degree is approximately $2\cdot\log n$ (around $12$ for $n=64$), which, in the bounded case, was exactly equal to the limit of $2\cdot h_{\max}=12$. Notably, the average degree remained tightly consistent across all algorithms. This confirms that the degree distribution of an SLN is heavily skewed, with the vast majority of nodes maintaining a low degree.
@@ 7 Related Work
Improved designs of overlay networks. Overlay networks have been the focus of computer scientists for a long while (see Lua et al. (2005) for a detailed survey). Many overlay networks of constant epected degree and logarithmic paths lengths are known, for example Naor and Wieder (2007) that extends the work of Fraigniaud and Gauron (2006), giving a general overlay construction such that for a network with degree $d$ and $n$ nodes, it is possible to guarantee path length of $\Theta(\log_{d}n)$. Lately, improving the design of networks given prior information about communication demand has attracted the attention of researchers. In particular, the work of Avin and Schmid (2018) set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking de Oliviera Souza et al. (2022); Pourdamghani et al. (2023); Figiel et al. (2025).
@@
Works closest to ours are by Avin et al. (2020c); Ciriani et al. (2007) in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input. This is particularly important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system. Also, the work of Mandal et al. (2015) discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide theoretical upper bounds for the cost of their overlay network. Lastly, we point out to other extensions of skip list, namely Skip Graph Aspnes and Shah (2003) and SkipNet Harvey et al. (2003) that, to facilitate overlay communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
@@
Learning-augmented algorithms. Given the rise of machine learning algorithms (which inherently comes with some errors), and after the pioneering work of Purohit et al. (2018) on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice. Such an approach has been used in various applications so far [^2] See a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching Lykouris and Vassilvitskii (2018), online TSP Gouleakis et al. (2023), list update Azar et al. (2025).
@@
Some of the more recent works focus on improving single-source search data structures. Initially, the works of Lin et al. (2022); Chen et al. (2025) provided learning-augmented search trees. The works that are closest to our work are by Fu et al. (2025); Zeynali et al. (2024), which studied learning-augmented single-source skip lists. In contrast, we consider a more challenging model, where the focus is on pairwise communications rather than communication from a single source. Recently, we became aware of a not peer-reviewed manuscript Aradhya and Scheideler (2025) that aims to augment overlay networks with learned advice; however, this paper looks into improving self-stabilization and recovery of overlay networks, not improving communication delay.